OpenStudy (anonymous):

integral (sqrt(z))/(1-z^3) dz

OpenStudy (anonymous):

\[\int\limits \frac{ \sqrt{z} }{ 1-z^3 }\]

OpenStudy (anonymous):

Is this what you meant?